OpenStudy (anonymous):

Evaluate each polynomial for x = -1 and y = 3. x + 3y

OpenStudy (anonymous):

x + 3y = (-1) + 3(3) = -1 + 9 = ??? i'm having trouble with the last step.. can you do it?